Find the absolute maximum and minimum values of f(x)=x−2x+4, if any, over the interval (−4,1].

absolute maximum is
and it occurs at x =

absolute minimum is
and it occurs at x
1
Expert's answer
2020-03-04T17:12:18-0500

Solution

Find the first derivative.

"f'(x) =2x-2"

Equate the first derivative to zero and solve the equation.

"2x-2=0"

"x=1"

The roots of the equation are critical points. Use the end points of the interval and all critical points on it to check for an absolute extremum in this interval.

"x=1,-4"

"f(1)=1^2-2*1+4=3"

"f(-4)=(-4)^2-(-4)*2+4=16+8+4=28"

We have that:

"f(1)<f(-4)"

It means that absolute maximum could be f(x) =28, but x=-4 is out of the domain (hence it is not attained), the absolute minimum is f(x) =3 and it occurs at x=1.


Answer:

the absolute minimum is "f(x) = 3" and it occurs at x=1, the absolute maximum is not reachable.